OpenStudy (rainbow_rocks03):

NO DIRECT ANSWERS. Choose the fraction that is equal to 1/4 A.3/4 B.3/12 C.4/1 D.4/4

OpenStudy (anonymous):

What do you think?...

OpenStudy (rainbow_rocks03):

I think A.

OpenStudy (rainbow_rocks03):

either A or B

OpenStudy (anonymous):

I'm pretty sure it's B

OpenStudy (cwrw238):

find a common factor of the numerator and denominator and divide them by it

OpenStudy (anonymous):

b cause 1*12 is 12 and 4*3 is 12 so they are equal

OpenStudy (anonymous):

Try 1/4 x 3/3 and then you have the answer.

OpenStudy (anonymous):

It is B here is why... 1/4= .25 3//24= .25

OpenStudy (rainbow_rocks03):

I did but it gives me the answer is decimal form

OpenStudy (anonymous):

0.25

OpenStudy (cwrw238):

there is only one fraction with a common factor (other than 1)

OpenStudy (rainbow_rocks03):

that is not an answer choice

OpenStudy (anonymous):

:P I am saying if the decimals are the same... The fractions will be also :)

OpenStudy (rainbow_rocks03):

it has to be in decimal form

OpenStudy (anonymous):

Try 1/4 x 3/3

OpenStudy (rainbow_rocks03):

I said i did and it gave me the answer in decimal form!

OpenStudy (anonymous):

The answer is B no matter what way you do it :P

OpenStudy (rainbow_rocks03):

ok thank you

OpenStudy (anonymous):

:)

OpenStudy (rainbow_rocks03):

when I open another question will someone help? bye

OpenStudy (cwrw238):

no direct answers please

OpenStudy (rainbow_rocks03):

that is what I said!

OpenStudy (cwrw238):

yes - but its hard to stop people doing it ...

OpenStudy (anonymous):

@cwrw238 It wasn't a direct answer :P She guessed B I explained why B was right but now everyone is confusing her...

OpenStudy (anonymous):

I was just letting her know not to let everyone confuse her by saying B is the answer no matter what way you do it...

OpenStudy (cwrw238):

@John3.16 - it was a direct answer - watch out for mods!

OpenStudy (anonymous):

No it wasn't :/ Not even close... I am not sure why you would even begin to consider it is I Direct answer. I have read the CoC a BUNCH and this is NOT a direct answer :P

OpenStudy (rainbow_rocks03):

JUST PLZ HELP ME ON THIS OTHER QUESTION PLZ AND PLZ STOP ARGUING.

OpenStudy (cwrw238):

whatever - but I know that people have been warned and suspended in similar circumstances

OpenStudy (rainbow_rocks03):

PLZ SOMEONE HELP ME!

OpenStudy (rainbow_rocks03):

if yall wan to chat private chat or go on a different site , I just need someone to help me plz!!!

OpenStudy (rainbow_rocks03):

SERIOUSLY YALL ARE ON THIS BOARD DOING NOTHING WHEN I AM ASKING FOR HELP!

OpenStudy (anonymous):

what do you need help with hun?

OpenStudy (anonymous):

.75 = 3/4= 1/4

OpenStudy (anonymous):

sorry... that was the last problem....